Audiocasette tapes cost 4 dollars each and compact discs cost 12 dollars each. If the number of audiocasette tape was 3 more than twice the number of compact discs and if a total of $92 was spent, how many audiocasette tapes and how many compact discs were purchased?
Answer by checkley71(8403)

4a+12c=92
a=2c+3
4(2c+3)+12c=92
8c+12+12c=92
20c=92-12
20c=80
c=80/20
c=4 compact disks were bought.
a=2*4+3
a=8+3
a=11 audio cassetts were bought.
proof
4*11+12*4=92
44+48=92
92=92